Find C(x,y) which is on 1/5 of the distance from A(6,8) to B(2,7)
Answer by Alan3354(69443)

Find C(x,y) which is on 1/5 of the distance from A(6,8) to B(2,7)
---------
Find x:
x = (2-6)/5 = -0.8
Find y:
y = (7-8)/5 = -0.2
---------
Add to point A
--> (5.2,7.8)